# Large Language Model Lateral Spear Phishing: A Comparative Study in Large-Scale Organizational Settings

### v1

@@ 3 Threat Model
The proliferation of LLMs marks a significant shift in the domain of lateral spear phishing, with these models introducing complexities of an unprecedented scale. This paper dissects the evolving threat landscape, pinpointing three pivotal roles: (1) Adversarial agents who harness compromised email accounts alongside LLMs to forge highly individualized and persuasive phishing communication. These communications are carefully tailored to target specific individuals within an organization [21], with the ultimate goal of illicitly obtaining sensitive information and breaching secure systems [22]. The spectrum of adversaries ranges from solitary hackers to state-sponsored groups [23], exploiting two key vulnerabilities. Firstly, there is the inherent trust employees place in communication that appears familiar. This trust allows adversaries to craft and insert phishing messages into an organization’s communication flow undetected, exploiting this implicit trust to their advantage [24, 25]. Secondly, the sophistication of LLMs presents a substantial challenge to traditional security systems, which are often not equipped to detect the nuanced and human-like text produced by such models, allowing this kind of content to slip past established detection mechanisms [26]. (2) Targets who receive these carefully engineered phishing communications. These targets span across all levels of hierarchy and profiles, and they may not always possess the necessary cybersecurity awareness to identify and avoid these sophisticated phishing attempts. (3) IT Professionals who are tasked with the development and administration of intricate email filters and security operation centers (SOCs) to combat and neutralize these evolving threats. They play a crucial role in implementing and managing advanced security solutions that can effectively detect and mitigate phishing. This study assumes that perpetrators have unauthorized access to both email accounts and Large Language Models (LLMs). These tools enable them to create highly personalized and convincing phishing messages that build trust and create urgency, often resulting in the unintended release of sensitive information or actions that undermine the security of the targeted organization. It is critical to reassess the basic principles of current cybersecurity measures. Typically, these measures focus on identifying well-known patterns of malicious activity, which may not suffice against the more sophisticated and subtly crafted attacks made possible by LLMs. These models can generate text that closely resembles genuine human writing. Furthermore, there is a common overestimation of the level of cybersecurity knowledge within an organization’s staff, ignoring the reality that awareness and training in cybersecurity vary widely among employees.
4 Lateral Phishing Simulation Infrastructure
4.1 Stakeholder Roles
4.1.1 Research Role
@@
For this study, we leveraged the already existing phishing awareness campaign infrastructure to test our research questions. We partnered with the organization’s cyber operations team and acted as advisors to administer the experiment. We helped the university’s cyber operations team prompt the LLM to develop effective phishing email templates. Researchers also acted as scientific advisors to guide the design of the experiments and research questions. Researchers only had access to anonymized data from the cyber operations team to perform analysis.
4.1.2 IT Cyber Operations Role
@@
Central to this cybersecurity framework is an overarching emphasis on training and awareness, especially against social engineering types of attacks. Monthly phishing simulation exercises cultivate an informed employee workforce and keep our cybersecurity insurance premiums low. Weekly communication regarding scams and red flags raises the institutions defense mechanisms against sophisticated cyber-threats. It is a constant balancing act of threat detection and response activities, systems patching and risk mitigation strategies complemented by phishing simulations and training.
4.1.3 Study Participants
Roughly 9000 employees were subject to the phishing exercises over the period of this 11 month phishing exercise, with 8995 employees receiving phishing emails in the last round of lateral phishing emails. During the last round of phishing emails, a total of 1345 employees held a supervisory role, which is important to note since several of the email templates used in this study relied on lateral phishing attacks where the email was addressed from a supervisor to a direct report, utilizing the trust and familiarity of a common type of organizational relationship.
4.2 Phishing Process
@@
4.2.1 Phishing Emails
This study includes the monthly phishing emails sent through the organization’s phishing training platform from January through November for a total of 11 months worth of phishing data with more than 20 total phishing templates used during this time. The phishing emails sent in the January through September were crafted without any feedback from researchers. The structure and topics of these emails varied, with some emails crafted to take advantage of recent news at the time such as student loan forgiveness. Some emails also included spoofed internal domains, and others included spoofed external domains such as Microsoft, Google, and Zoom. On the other hand, the phishing emails in October and November were crafted with the input of researchers. Like the January through September emails, October emails were also still human written, while the November emails were written using an LLM. The phishing templates used in October and November were crafted based on some of the observations of real-world social engineering tactics from the organization’s cyber operations team alongside researcher suggestions. As such, the phishing emails used in October and November were focused around a few specific themes. These phishing emails included emails from a supervisor to their direct report which leverage internal organizational information, and communications from other internal addresses that leverage external information.
The five phishing emails that were the primary focus of this study were the phishing emails in the October and November phishing exercises, with two different phishing templates occurring in October and three different phishing templates occurring in November. The two templates in October were human written. The first of these human written templates was designed to appear to be authored by the recipient’s supervisor and addressed the recipient by their first name, where the supervisor asks the recipient to click a link to upload some pictures for a presentation next week. The second human written template was designed to appear to be an urgent request from the organizaton’s Human Resources department addressing the recipient by their first name, asking the recipient to click a link to upload their address and phone number.
The next set of templates were the LLM authored phishing emails. There were three different LLM written email templates. The first two templates were written on the topic of an end-of-year holiday event, which appeared to be sent from the supervisor of the recipient, where the recipient was asked to write their suggestions to a spreadsheet that was linked in the email. The first LLM written template addressed the recipient by their first name, but the second LLM written template addressed the recipient by ”Hi Team”. GPT4 from ChatGPT was used to author these emails, being prompted to write an email template about gathering ideas from an end-of-year holiday event where the sender is the supervisor of the email recipient and is written in a mildly informal and friendly tone. The third LLM authored email template was a uniform template that was the same no matter the recipient and utilized external information about the organization. This template was written on the topic of new and exciting food options to be available at the university. The email was to appear to be written by an organizational account, but not someone who the recipients would know. GPT4 from ChatGPT was also used to author this email, and was asked to look at the university’s dining page to understand the style of the university dining writing and to write about new and exciting food options. In this email, the recipients were given a link to view the new menu and dining options.
4.2.2 Phishing Training and Warning
@@
4.2.3 Measurement
In conducting phishing exercises, we closely monitor employee responses, gathering essential metrics such as Sent, Opened, Clicked, Replied, Attachment, Data Entered (without saving credentials), and Reported. By aggregating these metrics with data from previous exercises, we formulate a risk score to evaluate the organization’s security posture and assess the susceptibility of our employee base to phishing attacks. Employees who enter data in these exercises are promptly enrolled in security awareness training, with a completion deadline of two weeks. Additionally, this risk score is benchmarked against industry data to provide insights, track historical trends, and gauge the maturity level of our cybersecurity posture.
@@ 4.2.3 Measurement
Month Email Recipients Emails Opened Link Clicked Data Entered Category of Phishing Emails
Nov. 2023 8,995 4,461 (49.59%) 1,501 (16.69%) 613 (6.81%)
Event Coordination Scam: Holiday Party Planning, New Dining Options
Oct. 2023 8,856 5,303 (59.88%) 2,352 (26.56%) 646 (7.29%)
Conference Pictures Request, Internal HR: Urgent Request
Sept. 2023 8,495 3,078 (36.23%) 398 (4.69%) 140 (1.65%)
Spoofed Suspicious Activity, Spoofed Domain Employee Verification
Aug. 2023 7,860 2,511 (31.95%) 181 (2.30%) 52 (0.66%)
Password Reset Fraud, Update Payment Information Scam
July 2023 6,750 3,084 (45.69%) 269 (3.99%) 100 (1.48%)
Internal Account Spoofed IT Support Login Info Required
June 2023 6,498 2,746 (42.26%) 262 (4.03%) N/A*
Microsoft Spoofed Domain without URL Link, Google Spoofed Domain without URL Link
May 2023 6,809 3,034 (44.56%) 843 (12.38%) 139 (2.04%)
Callback Voicemail Internal Account, Internal Account Retirement Plan Changes
Apr. 2023 6,759 3,620 (53.56%) 722 (10.68%) 53 (0.78%)
Internal HR: Have a message from HR, Message from Accounting
Mar. 2023 6,729 3,107 (46.17%) 769 (11.43%) 250 (3.72%)
Spoofed Domain: Need to Verify Information for W2, Personal Tax Information
Feb. 2023 6,614 2,776 (41.97%) 413 (6.24%) N/A*
Microsoft Spoofed Support for Password, Dropbox Spoofed Support for Password
Jan. 2023 6,271 2,334 (37.22%) 210 (3.35%) N/A*
Docusign Spoofed Document Need Signature, Student Loan Debt Relief Spoofed
TABLE II: Summary by month of results of phishing exercises conducted. Percentages shown are relative to the number of recipients. Also shown are short descriptions of some of the phishing emails sent during that month’s phishing exercise. Many emails were based on relevant news or events such as student loan forgiveness. Entries with N/A mean that this metric was not collected for that month’s phishing exercise.
4.3 Ethics, Safety and Privacy
@@
In this study, our team collaborated with the University’s cyber operations team, which had previously been conducting exercises using phishing emails to enhance awareness and educate users about the risks associated with such emails. The cyber operations team was responsible for managing all personally identifiable information (PII), ensuring its anonymization before providing the researchers with statistical data. Upon reviewing our research protocols, the IRB of our institution determined that our project did not fall under the regulatory criteria of the DHHS or the FFDA. Consequently, it did not require further oversight from the IRB.
5 Experiments and Discussion
5.1 RQ1: Effectiveness of Human-Crafted Lateral Phishing
To answer RQ1, we first show the results of 11 months of phishing exercises across a large educational organization. The results of these phishing exercises are presented in Table II, showing the number of phishing email recipients, phishing emails opened, phishing links clicked and data entered. Furthermore, a short description of the phishing emails in the month’s phishing exercise were included to provide some insight into what topics and themes were effective attacks. An analysis of the data reveals significant fluctuations in the success rates of these simulated phishing attacks. The effectiveness of lateral phishing can be primarily gauged by two key metrics: the percentage of recipients who opened the phishing link, and more critically, the percentage of those who further engaged by clicking on the link and entering their information.
The highest open rate was observed in October (59.88%), while the lowest was in August (31.95%). These variations can potentially be attributed to the nature of the email content, which may exploit the specific contexts and expectations within an educational setting. Emails with a sense of urgency or relevance to common educational or administrative themes, such as ”Internal HR: Urgent Request” or ”Conference pictures requested,” appeared to be more effective in prompting recipients to open the links.
The rate at which recipients clicked on the links and entered their information is a more direct measure of the phishing attack’s success. In this context, October again stands out with a click-through rate of 26.56% and an information entry rate of 7.29%. This suggests a significant vulnerability among the targeted individuals to such types of attacks.
5.2 RQ2: Impact of LLM-Crafted Lateral Phishing
@@
Email Email Template Email Recipients Emails Opened Link Clicked Data Entered
Human Written Supervisor to Direct Report 4,539 2,875 (63.34%) 1,701 (37.48%) 351 (7.73%)
LLM Generated Supervisor to Direct Report 6,070 3,479 (57.31%) 1,385 (22.81%) 607 (10.00%)
Human Written Critical Internal Communications 4,317 2,428 (56.24%) 651 (15.01%) 295 (6.83%)
LLM Generated Critical Internal Communications 2,925 982 (33.57%) 116 (3.97%) 6 (0.21%)
Total Human Written 8,856 5,303 (59.88%) 2,352 (26.56%) 646 (7.29%)
Total LLM Generated 8,995 4,461 (49.59%) 1,501 (16.69%) 613 (6.81%)
TABLE III: Human-crafted and LLM Crafted lateral phishing emails. Percentages are relative to the number of email recipients.
To next answer RQ2, Table III is shown for a comparative analysis of human-written versus LLM-generated lateral phishing emails. For the ”Supervisor to Direct Report” template, human-written emails exhibit a higher open rate (63.34%) compared to LLM-generated emails (57.31%). This suggests that human-written emails may initially engage recipients more effectively. However, in terms of eliciting specific responses such as clicking on links and entering data, LLM-generated emails demonstrate a noteworthy performance. The data entry rate for LLM-generated emails in this category is notably higher (10.00%) compared to human-written ones (7.73%), indicating their potential to effectively persuade recipients to take action.
In the case of ”Critical Internal Communications,” there is a marked difference in effectiveness between human-written and LLM-generated emails. The human-written emails outperform LLM-generated emails in all metrics. The lower impact of LLM-generated emails in this category can be attributed to the content’s relevance and urgency. The LLM-generated emails, focusing on less critical topics like dining options changes, failed to engage recipients as effectively as the human-written emails that dealt with more pressing matters like HR requests. This variation underscores the importance of content relevance in determining the effectiveness of phishing emails, whether human-crafted or AI-generated. The stark contrast in the effectiveness of these templates underscores the importance of context and personalization in phishing attacks. The success of the phishing emails from Table III aligns with insights from our discussions with the cyber operations team, which noted that timely attacks based on an organization’s calendar of events are effective in university-targeted phishing. This tendency further emphasizes the need for heightened vigilance and advanced detection methods in cybersecurity strategies, especially in anticipating and countering LLM-generated phishing emails that exploit specific organizational events and relationships.
The subtle differences in language and tone between the human-written and LLM-generated emails may contribute to these divergent outcomes. Despite these differences, the overall comparison across all categories indicates a close alignment in the effectiveness of human-written and LLM-generated emails. The total data entered rates are relatively similar (7.29% for human-written vs. 6.81% for LLM-generated), suggesting that under certain conditions, LLM-generated emails can achieve an impact comparable to that of human-written emails.
5.3 RQ3: LLM Phishing using Internal vs External Organization Information
@@
Email Template Email Recipients Emails Opened Link Clicked Data Entered
Targeted Internal Information 3,022 1,735 (57.41%) 606 (20.05%) 295 (9.76%)
Group Internal Information 3,048 1,744 (57.22%) 779 (25.56%) 312 (10.24%)
External Information 2,925 982 (33.57%) 116 (3.97%) 6 (0.21%)
TABLE IV: Internal vs External information for LLM Phishing. Percentage shown is percent relative to the number of email recipients.
To answer RQ3, we present Table IV. The analysis of LLM-generated spear phishing emails, distinguished by their use of internal versus external organizational information, provides critical insights into their varying effectiveness. The key distinction lies in how these emails engage with the recipients—emails categorized under ”Targeted Internal Information” utilize the recipient’s name, thereby creating a sense of direct and personal communication. This approach resulted in a 57.41% open rate, a 20.05% link click rate, and a 9.76% data entry rate, indicating a high level of effectiveness. The personalization aspect likely plays a significant role in this success, as recipients might perceive such emails as more credible and relevant.
In contrast, the ”Group Internal Information” emails, which address recipients with a general ”Hi Team,” also show strong engagement 57.22% open rate, 25.56% link click rate, and 10.24% data entry rate). Despite the lack of individual targeting, these emails still resonate effectively, possibly due to the perceived relevance to the group at large, which may prompt a sense of collective urgency or responsibility. However, when it comes to emails leveraging ”External Information,” such as discussing dining options at the university, the effectiveness markedly decreases. These emails exhibit only a 33.57% open rate, a 3.97% link click rate, and a notably low 0.21% data entry rate. This significant drop in engagement underscores the importance of contextual relevance and familiarity in spear phishing attacks. Emails with external content, lacking direct ties to internal operations or personalized touchpoints, are less likely to be perceived as legitimate or urgent, leading to lower interaction rates.
This comparison clearly demonstrates that LLM-generated spear phishing emails are more effective when they incorporate internal organizational information, whether targeted or group-oriented, as opposed to external content. The effectiveness of spear phishing is evidently heightened by the use of personalized or internally relevant information, emphasizing the need for heightened awareness and education about such sophisticated phishing tactics within organizations.
5.4 RQ4: Effectiveness of Phishing Training and Warnings
To answer RQ4, we perform a sub-analysis that focuses on a specific subset of individuals: those who previously entered data in the human-crafted phishing email experiment during the October phishing exercises and subsequently received phishing training. This analysis is crucial in understanding the effectiveness of training in mitigating susceptibility to phishing, particularly in the context of sophisticated LLM-crafted lateral phishing emails.
The results from this sub-analysis, as shown in Table V, reveal some concerning trends. Among those who received phishing training, a high percentage still fell prey to the LLM-crafted phishing emails, especially in the supervisor-to-direct-report scenarios. In the first scenario, where the supervisor’s email mentioned the recipient’s name, 77.88% of the 208 targeted individuals opened the email, 36.05% clicked on the link, and a significant 24.51% entered data. Similarly, in the second scenario, with the team mentioned, 81.53% of 222 targeted individuals opened the email, 40.09% clicked on the link, and 20.72% entered data. These rates are noticeably higher than the general population’s response in the previous experiments, indicating that even after training, these individuals were 2-3 times more vulnerable to entering data from LLM-crafted lateral phishing attacks than the rest of the population. This points to the need for more effective training measures to reduce phishing vulnerability.

### v2

@@ IV Threat Model
The proliferation of LLMs marks a significant shift in the domain of lateral phishing, with these models introducing complexities of an unprecedented scale. This paper considers two roles: (1) Adversarial agents who harness compromised email accounts alongside LLMs to forge highly individualized and persuasive phishing communication. These communications are carefully tailored to target specific individuals within an organization [40], with the ultimate goal of illicitly obtaining sensitive information and breaching secure systems [41]. The spectrum of adversaries ranges from solitary hackers to state-sponsored groups [42]. There is inherent trust that employees place in communication that appears familiar. This trust allows adversaries to craft and insert phishing messages into an organization’s communication flow undetected, exploiting this implicit trust to their advantage [43, 44]. (2) Targets who receive these carefully engineered phishing communications. These targets span across all levels of hierarchy and profiles, and they may not always possess the necessary cybersecurity awareness to identify and avoid these sophisticated phishing attempts. This study assumes that perpetrators have unauthorized access to both email accounts and LLMs. These tools enable them to create highly personalized and convincing phishing messages that utilize trust and create urgency, often resulting in the unintended release of sensitive information or actions that undermine the security of the targeted organization. LLMs can generate text that closely resembles genuine human writing. Furthermore, there is a common overestimation of the level of cybersecurity knowledge within an organization’s staff, ignoring the reality that awareness and training in cybersecurity vary widely among employees.
V Lateral Phishing Simulation Infrastructure
Figure 2 illustrates the setup of the red-teaming infrastructure used for lateral phishing simulations. The infrastructure consists of several key components working in concert. For the purposes of the red-teaming exercise, the simulated phishing emails bypass organizational email filters, mimicking either a compromised employee account or an organizational email address. The cyber operations team leverages their phishing simulation dashboard to incorporate organizational context such as employee names into the phishing emails. Additionally, in the context of this study, organization news from public sites were used to identify relevant upcoming events which would be the subject of one of the types phishing emails we studied. These simulated phishing emails are then distributed to targeted employee mailboxes. The cyber operations team records all employee interactions with the phishing emails, from whether the email was opened, to the link having been clicked, to having input login credentials into the phishing link. If an employee enters their login credentials through the phishing link, they are immediately redirected to a warning page (shown in Figure 3) that informs them they have participated in a simulated phishing test.
V-A Stakeholder Roles
V-A1 Research Role
@@
For this study, we leveraged the already existing phishing awareness campaign infrastructure to test our research questions. We partnered with the organization’s cyber operations team and acted as advisors to administer the experiment. We worked with the university’s cyber operations team prompt the LLM and human to develop effective phishing email templates. Researchers also acted as scientific advisors to guide the design of the experiments and research questions. Researchers only had access to anonymized data from the cyber operations team to perform analysis.
V-A2 IT Cyber Operations Role
@@
Central to this cybersecurity framework is an overarching emphasis on training and awareness, especially against social engineering types of attacks. Monthly phishing simulation exercises cultivate an informed employee workforce and keep cybersecurity insurance premiums low. Weekly communication regarding scams and red flags raises the institutions defense mechanisms against sophisticated cyber-threats. It is a constant balancing act of threat detection and response activities, systems patching and risk mitigation strategies complemented by phishing simulations and training.
V-A3 Study Participants
In this study, 9129 employees were subjected to the phishing exercise. Among these employees, 1386 held supervisory roles. This distinction is important because one of the email templates used in the study relied on lateral phishing attacks, where the email was addressed from a supervisor to a direct report, exploiting the trust and familiarity inherent in this common type of organizational relationship. We also analyzed the results of the phishing experiment by department and role. The organization had 302 unique departments, which were manually mapped to 10 department categorizations. The study encompassed a range of employee types across the university. There were 1,219 unique job roles, which were also manually mapped to 9 job role categorizations.
V-B Phishing Process
@@
V-C Phishing Emails
To answer our research questions, we devised a standardized methodology for formulating the email templates. Both the human participant, a professional with a background in communications representing the upper-end of human ability in crafting persuasive and coherent messages, and the LLM, specifically ChatGPT GPT-4, were given identical prompts to generate phishing emails. To maintain consistency and fairness in the evaluation, the initial drafts produced by both the human and the LLM were subjected to a review process by the research team. This process involved making minimal adjustments to the human-composed email to align it closely with its LLM counterpart in terms of the sender’s identity, the email’s subject matter, the degree of personalization, and the conveyed sense of urgency. Utmost care was taken to ensure that the writing style of each email was maintained, preserving the unique characteristics of human and LLM-generated content. This approach was carefully chosen to ensure that any observed variances in the effectiveness of the phishing attempts could be attributed primarily to the intrinsic differences between human and LLM-generated content, rather than extraneous factors. Across all emails, a sense of urgency was required as part of the prompt to generate the emails, as this is a common tactic employed in phishing attempts to pressure recipients into taking immediate action without thoroughly considering the consequences. We tested three types of lateral phishing: emails from a supervisor, organizational-level emails, and emails from another internal employee who the recipient does not know, to assess the effectiveness of different attack vectors and identify potential vulnerabilities within the institution’s hierarchy and communication channels. For all phishing email templates generated, the recipient is asked to complete some action by clicking a link in the email so that we could record the recipient’s vulnerability to clicking a phishing link and entering data or credentials in the web interface the link leads to. In the following subsubsections, we outline the prompts given to the LLM and human to generate the lateral phishing email templates and reasoning behind why these topics were chosen.
V-C1 Supervisor to direct report
Prompt: ”Write a generic email from a supervisor to direct report, where the direct report is instructed to click some link with some urgency to enter some information about professional development.”
In this scenario, we asked both the LLM and the human to compose an email from a supervisor to their direct report, using the prompt above. The email instructs the recipient to click on a link with urgency to enter information about professional development. This topic was chosen because it represents a plausible situation in which a supervisor might make such a request of their direct report. The inherent power dynamics and trust within the supervisor-direct report relationship could potentially make this type of phishing attempt highly effective, as employees may feel compelled to comply with their supervisor’s instructions without questioning the legitimacy of the request. In subsequent sections we describe this email topic as ”Supervisor to Direct Report”.
V-C2 Organizational-level email on a timely event
Prompt: ”Write a generic email on some timely internal communications where the sender would be from some organizational level email address, where the email asks the recipient to enter some information in a link with some urgency. The topic of the email should about making reservations for eyewear relating to the upcoming solar eclipse.”
For this topic, the LLM and human were tasked with writing a generic email using the prompt above. The email is sent from an organizational-level email address, asking the recipient to enter information via a link with urgency. The specific subject of the email was making reservations for eyewear related to an upcoming solar eclipse. This topic was selected because the solar eclipse was a well-known event that was going to occur, and the institution was known to be hosting events related to the eclipse. Moreover, the organization had previously experienced phishing attacks that capitalized on timely events, making this a relevant and realistic scenario to test the effectiveness of LLM and human-generated phishing emails in exploiting current events and institutional context. In subsequent sections we describe this email topic as ”Timely Phishing”.
V-C3 Common Phishing Email from Internal Employee Account
Prompt: ”Write a generic email about extending a job offer where the recipient would be doing some remote work and is asked to click on a link to enter some information with some urgency.”
In this case, both the LLM and human were asked to compose a generic email using the prompt above. The email is about extending a job offer, where the recipient would be doing some remote work and is asked to click on a link to enter information with urgency. According to the organization’s cyber operations team, job offer phishing schemes are the most common type of phishing attack faced by the institution. By using this topic, we can assess the effectiveness of LLM and human-generated phishing emails in replicating a prevalent real-world threat. Additionally, the scenario involves an email from an unfamiliar internal employee, testing the recipient’s ability to discern the legitimacy of the request without the inherent trust that may be present in established relationships within the organization. In subsequent sections we describe this email topic as ”Common Phishing”.
V-D Phishing Training and Warning
@@
Email Author Email Topic Email Recipients Emails Opened Link Clicked Data Entered
Human Supervisor to Direct Report 1541 955 (61.97%) 334 (21.67%) 164 (10.64%)
LLM Supervisor to Direct Report 1490 919 (61.68%) 318 (21.34%) 165 (11.07%)
Human Timely Phishing 1591 574 (36.08%) 106 (6.66%) 40 (2.51%)
LLM Timely Phishing 1479 597 (40.37%) 153 (10.34%) 62 (4.19%)
Human Common Phishing 1520 831 (54.67%) 68 (4.47%) 18 (1.18%)
LLM Common Phishing 1508 715 (47.41%) 73 (4.84%) 14 (0.93%)
Total Human Written 4652 2360 (50.73%) 508 (10.92%) 222 (4.77%)
Total LLM Generated 4477 2231 (49.83%) 544 (12.15%) 241 (5.38%)
Total 9129 4591 (50.29%) 1052 (11.52%) 463 (5.07%)
TABLE II: Comparison of human-crafted and LLM-generated phishing email effectiveness. Percentages are relative to the number of email recipients.
V-E Measurement
In conducting phishing exercises, we closely monitor employee responses, gathering essential metrics such as Sent, Opened, Clicked, Replied, Attachment, Data Entered (without saving credentials), and Reported. By aggregating these metrics with data from previous exercises, we formulate a risk score to evaluate the organization’s security posture and assess the susceptibility of our employee base to phishing attacks. Employees who enter data in these exercises are promptly enrolled in security awareness training, with a completion deadline of two weeks. Additionally, this risk score is benchmarked against industry data to provide insights, track historical trends, and gauge the maturity level of the organization’s cybersecurity posture.
@@ V-E Measurement
V-F Ethics, Safety and Privacy
@@
In this study, our team collaborated with the University’s cyber operations team, which had previously been conducting exercises using phishing emails to enhance awareness and educate users about the risks associated with such emails. The cyber operations team was responsible for managing all personally identifiable information (PII), ensuring its anonymization before providing the researchers with statistical data and other qualitative insights. Upon reviewing our research protocols, the IRB of our institution determined that our project did not fall under the regulatory criteria of the DHHS or the FFDA. Consequently, it did not require further oversight from the IRB. While there could be concerns about the lack of informed consent, these phishing simulation exercises are currently already conducted throughout the organization on a regular basis as a part of standard organizational security practices, with researchers simply helping to inform the methodology to generate emails. The cyber operations team communicated the details of the proposed lateral phishing exercise with IT governance in the University system to obtain approval to conduct the exercise, where the results of the exercise could help inform the organization quantify the potential threat of LLMs towards phishing within the organization and potentially inform security practices and training in the future. To reduce the psychological impact, the cyber operations team established a trusted agents committee that meets prior to the launch of phishing campaigns. This committee convenes before phishing campaigns and engages in a discussion forum where members review and approve the email templates, offer feedback, and adjust accordingly. The committee consists of more than eight members representing different areas including chief of staff, risk management and legal affairs, College of Business, Academic affairs operations, Human Resources, strategic enrollment and student affairs, and research integrity, in order to limit potential harms regarding the phishing campaign.
VI Experiments and Discussion
VI-A RQ1: Human vs LLM Lateral Phishing emails
@@
Email Author Email Topic Email Recipients Emails Opened Link Clicked Data Entered
Human Supervisor to Direct Report 249 178 (71.49%) 47 (18.88%) 12 (4.82%)
LLM Supervisor to Direct Report 241 176 (73.03%) 46 (19.09%) 20 (8.30%)
Human Timely Phishing 225 120 (53.33%) 22 (9.78%) 12 (5.33%)
LLM Timely Phishing 227 150 (66.08%) 40 (17.62%) 15 (6.61%)
Human Common Phishing 228 159 (69.74%) 2 (0.88%) 1 (0.44%)
LLM Common Phishing 216 126 (58.33%) 1 (0.46%) 0 (0.00%)
Total Human Written 702 457 (65.10%) 71 (10.11%) 25 (3.56%)
Total LLM Generated 684 452 (66.08%) 87 (12.72%) 35 (5.12%)
Total 1386 909 (65.58%) 158 (11.40%) 60 (4.33%)
TABLE III: Comparison of human-crafted and LLM-generated phishing email effectiveness. Percentages are relative to the number of email recipients. Filtered down to only those who supervise at least one other employee the organization.
open click enter
supervisor to direct report 0.167071 0.222355 -0.38169
timely phishing -2.44353* -3.66803* -2.59189*
common phishing 3.994332* -0.47943 0.688341
all 0.85823 -1.8414* -1.3299
TABLE IV: t-tests comparing human vs LLM emails across lateral phishing topics. Positive values indicate greater effect by human-authored emails; negative values by LLM-authored emails. Significant differences (*) are at p ¡ 0.05 after False Discovery Rate adjustment.
To answer Research Question 1, which compares the effectiveness of lateral phishing emails generated by LLMs and those crafted by humans within an educational organization, we analyze the data presented in Table II and the t-test results for statistical significance of this comparison in Table IV.
@@
Department Email Recipients Emails Opened Link Clicked Data Entered
College of Business 603 388 (64.34%) 102 (16.92%) 37 (6.14%)
College of Education and Human Development 480 252 (52.50%) 45 (9.38%) 23 (4.79%)
College of Engineering 638 340 (53.29%) 81 (12.70%) 29 (4.55%)
College of Liberal and Fine Arts 743 429 (57.74%) 87 (11.71%) 46 (6.19%)
College of Sciences 1270 768 (60.47%) 155 (12.20%) 57 (4.49%)
College of Architecture, Construction and Planning 166 88 (53.01%) 22 (13.25%) 12 (7.23%)
Honors College 509 288 (56.58%) 74 (14.54%) 36 (7.07%)
University College 315 168 (53.33%) 40 (12.70%) 15 (4.76%)
Administration 2685 1214 (45.21%) 300 (11.17%) 139 (5.18%)
Other 1720 656 (38.14%) 146 (8.49%) 69 (4.01%)
Total 9129 4591 (50.29%) 1052 (11.52%) 463 (5.07%)
TABLE V: Effectiveness of phishing emails across different departments. Percentages are relative to the number of email recipients
Table II provides an overview of the phishing email performance metrics for both human-crafted and LLM-generated emails across three different lateral phishing email topics: supervisor to direct report, timely phishing, and common phishing. The data shows that the overall percentage of emails opened, links clicked, and data entered are relatively similar between human-written and LLM-generated emails. For the ”supervisor to direct report” scenario, the human-crafted email had a slightly higher percentage of emails opened (61.97% vs. 61.68%) and links clicked (21.67% vs. 21.34%), while the LLM-generated email had a slightly higher percentage of data entered (11.07% vs. 10.64%). In the ”timely phishing” scenario, the LLM-generated email outperformed the human-crafted email across all metrics, with a higher percentage of emails opened (40.37% vs. 36.08%), links clicked (10.34% vs. 6.66%), and data entered (4.19% vs. 2.51%). For the ”common phishing” scenario, the human-crafted email had a higher percentage of emails opened (54.67% vs. 47.41%), while the LLM-generated email had a slightly higher percentage of links clicked (4.84% vs. 4.47%) and a lower percentage of data entered (0.93% vs. 1.18%).
To determine the statistical significance of these differences, we conducted t-tests comparing the human-crafted and LLM-generated emails for each phishing email topic, shown in Table IV. Because performing multiple comparisons increases the risk of false positives, we applied the Benjamini-Hochberg False Discovery Rate (FDR) correction. This correction adjusts the p-values to control the proportion of type I errors (i.e., falsely identifying a difference as significant). Consequently, only comparisons with adjusted p-values less than 0.05 are considered statistically significant, meaning there is less than a 5% probability that these observed differences occurred by chance. Specifically, for the ’timely phishing’ scenario, the LLM-generated email performed significantly better than the human-crafted email in terms of emails opened (t = -2.44, p <<< 0.05), links clicked (t = -3.67, p <<< 0.05), and data entered (t = -2.59, p <<< 0.05). In the ’common phishing’ scenario, the human-crafted email had a significantly higher percentage of emails opened (t = 3.99, p <<< 0.05), while no significant differences were observed for links clicked or data entered. For the ’supervisor to direct report’ scenario, no statistically significant differences were found. These t-values indicate both the direction and magnitude of the differences, where larger absolute values correspond to stronger effects. The p <<< 0.05 threshold confirms that the chance of these findings occurring randomly is below 5%.
These findings suggest that LLM-generated phishing emails can be as effective as, and in some cases even more effective than, human-crafted phishing emails. The comparable performance of LLM-generated emails in the ”supervisor to direct report” scenario highlights the potential threat of LLM phishing attacks, as they can exploit the trust and power dynamics within an organization. In conclusion, our analysis reveals that LLM-generated lateral phishing emails can be as effective as those crafted by humans within an educational organization. This finding underscores the need for organizations to adapt their cybersecurity strategies to address the growing threat of AI-powered phishing attacks and to invest in employee training and awareness programs that specifically target the unique characteristics of LLM-generated phishing emails.
Given the effectiveness of the lateral phishing emails from the supervisor to direct report, we further investigated the susceptibility of those who hold supervisory roles themselves since a single compromised account where they hold a supervisory role could lead to an effective propagation of an attack. Comparing the results of the phishing experiment for supervisors (Table III) to the entire organization (Table II), we observe that supervisors have significantly higher email open rates across all email topics. The most notable difference is in the ”Timely Phishing” category, where supervisors have an open rate of 59.71% compared to 38.23% for the entire organization. However, when examining link click rates and data entry rates, supervisors show varying levels of susceptibility depending on the email topic. For the ”Supervisor to Direct Report” category, supervisors have lower link click (18.98%) and data entry (6.56%) rates compared to the entire organization (21.51% and 10.86%, respectively). This suggests that supervisors may be more cautious when receiving emails appearing to be from their own supervisors. In the ”Timely Phishing” category, supervisors have higher link click (13.72%) and data entry (5.97%) rates compared to the entire organization (8.44% and 3.32%, respectively). This indicates that supervisors may be more vulnerable to timely and relevant phishing attempts.
VI-B RQ2: Breakdown by job role and department affiliation